what is the summation notation of 1000,100,10,1,0.1,0.01

Found 2 solutions by Edwin McCravy, rothauserc:
Answer by Edwin McCravy(20056)
Answer by rothauserc(4718)

This is a geometric progression with common ratio = 0.10 and a(1) = 1000
:
S(n) = a(1) * (1 - r^n) / (1 - r)
:
S(6) = 1000 * (1 - 0.10^6) / (1 - 0.10) = 1111.11
: